In the lab, Kevin has two solutions that contain alcohol and is mixing them with each other. He uses 3 times as much Solution A as Solution B. Solution A is 18% alcohol and Solution B is 11% alcohol. How many milliliters of Solution B does he use, if the resulting mixture has 390 milliliters of pure alcohol?
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)
Answer by ikleyn(52781)

.
In the lab, Kevin has two solutions that contain alcohol and is mixing them with each other.
He uses 3 times as much Solution A as Solution B.
Solution A is 18% alcohol and Solution B is 11% alcohol.
How many milliliters of Solution B does he use, if the resulting mixture has 390 milliliters
of pure alcohol?
~~~~~~~~~~~~

x mL of solution B  and  3x mL of solution A.

0.11x mL of the pure alcohol goes with x mL of solution B,

and 0.18*(3x) mL of the pure alcohol goes with x mL of solution A.


Write an equation for the pure alcohol volume

    0.11x + 0.18*(3x) = 390  mL


Simplify and find x

    0.11x + 0.54x = 390

         0.65x    = 390

             x    = 390/0.65 = 600.

ANSWER.  600 mL of solution B should be used.

Solved.

-----------------

It is a standard and typical mixture problem.

For introductory lessons covering various types of mixture word problems see
    - Mixture problems
    - More Mixture problems
    - Solving typical word problems on mixtures for solutions
    - Typical word problems on mixtures from the archive
in this site.

You will find there ALL TYPICAL mixture problems with different methods of solutions,
explained at different levels of detalization,  from very detailed to very short.

Read them and become an expert in solution mixture word problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this textbook in the section "Word problems" under the topic "Mixture problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.